OpenStudy (anonymous):

Related Rates? A water tank is shaped as an inverted cone with a heigh 10 ft and the radius of the circle at the top 10 ft. Volume of water drains from bottom at rate proportional to height of water level.

OpenStudy (anonymous):

So \[\frac{ dV }{ dt }=c \times h(t)\] where V(t) is water in the tank, t is time, and c is constant, and h(t) is height of water at time t. When the tank is full, the drain is opened and measurement shows that water level drops at instantaneous rate of 1 ft/hr. What is the instantaneous rate of change for the volume of the water when the drain is first opened?

OpenStudy (anonymous):

\[V=\frac{ \pi r^2h }{ 3 }\]

OpenStudy (anonymous):

 At any given time, the \(r\) leg of the triangle will be the same length as the \(h\) leg, so \[V=\frac{\pi r^2h}{3}=\frac{\pi r^3}{3}=\frac{\pi h^3}{3}\] Let's use the volume expression as a function of the water level \(h\). At time \(t=0\), you know that the rate of change of the water level is \(\dfrac{dh}{dt}=-1\) (negative because the tank is being drained). Differentiating the volume expression with respect to time gives \[\frac{dV}{dt}=\pi h^2\frac{dh}{dt}=-\pi h^2\] Also, at time \(t=0\), since the tank is full, you know that \(h=10\). Simple computation from here. Oh, and keep track of your units.